### v2

# Teaching Large Language Models to Reason through Learning and Forgetting

Abstract
Leveraging inference-time search in large language models has proven effective in further enhancing a trained model’s capability to solve complex mathematical and reasoning problems. However, this approach significantly increases computational costs and inference time, as the model must generate and evaluate multiple candidate solutions to identify a viable reasoning path. To address this, we propose an effective approach that integrates search capabilities directly into the model by fine-tuning it using both successful (learning) and failed reasoning paths (forgetting) derived from diverse search methods. While fine-tuning the model with these data might seem straightforward, we identify a critical issue: the model’s search capability tends to degrade rapidly if fine-tuning is performed naively. We show that this degradation can be substantially mitigated by employing a smaller learning rate. Extensive experiments on the challenging Game-of-24 and Countdown mathematical reasoning benchmarks show that our approach not only outperforms both standard fine-tuning and inference-time search baselines but also significantly reduces inference time by 180.111Code is available at https://github.com/twni2016/llm-reasoning-uft.
1 Introduction
The last few years has seen the rapid development of large language models (LLMs) (Vaswani et al., 2017; Achiam et al., 2023) and their applications to a diverse set of tasks. In particular, these LLMs have been tested on challenging benchmarks requiring high-level reasoning across domains such as mathematics (Glazer et al., 2024), abstract reasoning (Chollet et al., 2024), code generation (Zhuo et al., 2024), and science (Mitchener et al., 2025).
Chain-of-Thought (CoT) and inference-time search (search) have been proposed to enhance LLM generalization to reasoning problems at test-time. Specifically, given a task description and input , CoT (Wei et al., 2022) outputs a single reasoning path via greedy decoding, or samples multiple paths to compute average performance. In contrast, search-based methods (Lightman et al., 2023; Yao et al., 2023a; Snell et al., 2024) explicitly generate intermediate reasoning (possibly with multiple passes) before extracting via a reward model or self-evaluation. Despite inference-time search’s effectiveness, it comes with greater computational costs (e.g., Fig. 1 and Chen et al. (2024b; a)).
A beneficial outcome of utilizing inference-time sampling strategies, such as CoT and search-based methods, is the generation of rich datasets. These datasets can subsequently be fed back into the model for further fine-tuning, potentially enhancing the base model’s overall performance. Prior work has primarily focused on distilling the knowledge from successful reasoning paths, drawing from CoT-generated data (Zelikman et al., 2022; Yuan et al., 2023; Singh et al., 2023), paths from search trees (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a), or full search traces (Gandhi et al., 2024; Lehnert et al., 2024). However, in many complex problems, failed (reasoning) paths far outnumber successful ones, representing an underutilized resource that can be effectively used for training. Additionally, diverse search algorithms yield varying solutions, potentially improving data coverage and reasoning quality. Nonetheless, effectively leveraging both successful and failed reasoning paths and integrating data from both CoT and search methods to enhance reasoning remains largely underexplored.
In this paper, we propose a straightforward and simple pipeline consisting of three phases designed to enhance the reasoning capabilities of large language models: data generation, fine-tuning, and evaluation. First, we collect heterogeneous reasoning data from both CoT and the search-based methods, and convert the data into a unified CoT-style path format. Next, we fine-tune the base LLM by learning successful paths with the supervised fine-tuning (SFT) and forgetting failed paths with the unlikelihood loss. We refer to this approach as unlikelihood fine-tuning (UFT). Finally, we evaluate the fine-tuned model on test sets using both CoT and inference-time search, with a primary focus on efficient CoT inference. We demonstrate the effectiveness of UFT on two math reasoning domains: Game-of-24 and Countdown, using Qwen2.5-Math 1.5B and 7B (Yang et al., 2024). We summarize our key findings below:
1.
Data quality drives performance: On Countdown, replacing CoT data with high-quality classic search data boosts CoT performance from to , significantly outperforming the inference-time search baseline () with a reduction in inference cost.
2.
Forgetting improves performance: Incorporating failed reasoning paths using UFT improves CoT performance by – on average, with up to gains in best-case scenarios.
3.
Controlled learning rate helps maintain search capability: Naive fine-tuning can quickly degrade a model’s search ability. A substantially reduced learning rate helps prevent this, making it a simple but crucial step when fine-tuning with CoT-style data.
2 Related Work
LLM inference-time search.
Inference-time search has shown remarkable success across reasoning (Wang et al., 2022; Stechly et al., 2024; Hao et al., 2024; Snell et al., 2024) and planning tasks (Valmeekam et al., 2023; Zheng et al., 2024; Bohnet et al., 2024). Approaches include best-of-n sampling (Nakano et al., 2021; Wang et al., 2022), multi-agent debates (Du et al., 2023), and iterative correction via predefined rules (Bai et al., 2022) or self-refinement (Shinn et al., 2023; Madaan et al., 2024). A key direction, which this paper focuses on, integrates classic search algorithms with LLMs, employing BFS (Yao et al., 2023a; Xie et al., 2023), DFS (Qin et al., 2023), A* (Zhuang et al., 2023; Meng et al., 2024; Koh et al., 2024) and MCTS (Hao et al., 2023; Zhou et al., 2023; Zhao et al., 2024; Xie et al., 2024; Gao et al., 2024). While effective, inference-time search methods remain computationally expensive: self-refinement requires multiple expensive passes, multi-agent approaches demand substantial GPU memory, and search trees grow exponentially with depth (Chen et al., 2024b).
LLM policy distillation from synthetic data.
Synthetic data, generated by classic algorithms or LLMs, is widely used for SFT in reasoning tasks to mitigate the scarcity of human-annotated datasets (Liu et al., 2024). However, despite its abundance, synthetic responses may fail to solve the task. To address this, most prior work distill reasoning knowledge to policy (Hinton et al., 2015) by selecting only correct responses. Policy distillation strategies can be categorized based on whether the source and target data originate from CoT or search algorithms. CoT-to-CoT distillation learns CoT reasoning from correct CoT paths filtered by rewards, widely used in practice (Zelikman et al., 2022; 2024; Uesato et al., 2022; Yuan et al., 2023; Gulcehre et al., 2023; Singh et al., 2023). Search-to-search distillation either learns a meta-policy by imitating entire search traces (Yang et al., 2022; Gandhi et al., 2024; Lehnert et al., 2024) or refines the proposal policy in tree search by imitating extracted paths (Feng et al., 2023; Tian et al., 2024; Zhang et al., 2024a). Notably, CPO (Zhang et al., 2024c) performs search-to-CoT distillation using preference pairs from Tree-of-Thought without a verifier. In comparison, we focus on CoT, search-to-CoT distillation from diverse reasoners and distill unpaired positive and negative data to the policy, labeled by a rule-based verifier.
Offline fine-tuning with negative data.
The abundance of failed responses compared to successful ones provides an opportunity to leverage them for more efficient and generalizable learning. In case of LLM-based reasoning, a common approach is to train a reward model (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023; Feng et al., 2023; Zhang et al., 2024a; Hosseini et al., 2024) for response re-ranking and search during inference. However, this does not inherently improve the base model as a policy. Another approach extends vanilla DPO (Rafailov et al., 2023) by building preference pairs and fine-tuning the policy with offline preference optimization (Pang et al., 2024; Zhang et al., 2024c; Pal et al., 2024; Setlur et al., 2024). Beyond reasoning, LLM safety literature explores methods for unlearning harmful text (Welleck et al., 2019; Keskar et al., 2019; Yao et al., 2023b; Zhang et al., 2024b). Instead of forming preference pairs or incurring computational overhead to identify high-credit steps, we adopt unlikelihood training (Welleck et al., 2019) to directly forget failed responses as an auxiliary loss, avoiding the limitation of optimizing only relative probabilities within preference pairs (Tuan & Wang, 2024; Feng et al., 2024).
3 Preliminaries
Reasoning task as an MDP.
We formalize a reasoning task as a token-level Markov decision process (MDP) . The initial state consists of a tokenized sequence representing the task description and input. An action is chosen based on the current state , which is the sequence of all previous tokens: . A terminal state is reached upon generating a special end-of-sequence token. The ground-truth reward function is a rule-based verifier that checks the correctness of , yielding , where means failure and means success. The goal is to find a policy that generates an action sequence (denoted as ) which maximizes the reward: . We refer to the terminal state as a (reasoning) path.
LLM reasoners.
We use an LLM policy , parameterized by , along with a search algorithm as the reasoner. We denote this LLM reasoner as and focus on three popular LLM reasoners (Hao et al., 2024): (1) CoT (Wei et al., 2022) that uses greedy search, (2) Tree-of-Thought (ToT) (Yao et al., 2023a) that uses beam search, (3) Reasoning-via-Planning (RAP) (Hao et al., 2023) that uses MCTS (Kocsis & Szepesvári, 2006).
CoT directly outputs one (via greedy decoding) or multiple reasoning paths (under non-zero temperature) without intermediate search process222CoT’s performance is often measured by the average success rate over generated reasoning paths, also known as the pass@1 metric (Chen et al., 2021; Guo et al., 2025), as used in this work.. On the contrary, ToT and RAP are inference-time search methods, first constructing a search tree of reasoning paths before selecting a final path as the answer. Their performance is evaluated based on the success rate of the selected reasoning path. Please see Appendix A for more details.
4 Method: Unlikelihood Fine-Tuning for Reasoning
In this section, we describe our approach to solving a reasoning task, as outlined in Fig. 2. Our approach involves three stages: (1) generating reasoning data, (2) fine-tuning the base LLM from unpaired correct and failed paths and (3) evaluating it with reasoning algorithms.
4.1 Data Generation Using LLM and Classic Reasoners
For a given reasoning task, we generate synthetic reasoning data from multiple reasoners. LLM reasoners share a common base policy but differ in their search algorithms , which include CoT and inference-time search methods such as ToT and RAP. Optionally, reasoning paths can also be generated using classical search algorithms (classic reasoners) such as DFS and BFS, which rely on an external verifier rather than LLMs.
For search-based methods, we extract all root-to-leaf paths from the search tree – rather than only the final selected path – to follow the format of CoT-style paths. This enables (1) unifying heterogeneous search traces into a common format to improve coverage and data quality, and (2) efficient inference, as CoT inference is fastest. We aggregate CoT-style paths from all reasoners to compose the training dataset :
where is the set of considered reasoners, and is the set of all reasoning paths produced by reasoner . Note that classic reasoners do not depend on . Each path is labeled by a verifier as correct or failed , splitting into a correct dataset and a failed dataset .
4.2 Unlikelihood Fine-Tuning from Unpaired Correct and Failed Paths
While prior work typically fine-tune LLMs on correct CoT-generated paths (Yuan et al., 2023; Singh et al., 2023) or on preference pairs of correct and failed responses (Pang et al., 2024; Zhang et al., 2024c), we extend these approaches in two directions: (1) incorporating diverse reasoners beyond CoT to augment the training data, (2) fine-tuning the model to avoid and forget failed reasoning patterns using only failed examples. Unlike preference-based methods, our approach scales better by leveraging unpaired negative data, without requiring success-failure pairs for the same question.
Learning to follow correct reasoning paths.
First, we fine-tune the model using negative log-likelihood (NLL) loss – also known as supervised fine-tuning (SFT) (Ouyang et al., 2022; Cen et al., 2025) – on correct reasoning paths collected from multiple reasoners:
Examining the gradient of the NLL objective (Eq. 2), , reveals its connection to REINFORCE (Williams, 1992) with a binary (indicator) reward function, as noted in prior work (Zelikman et al., 2022; Gulcehre et al., 2023; Singh et al., 2023)333This resemblance holds when is close to on-policy data (Fakoor et al., 2020), e.g., CoT-generated, but may break down as model weights shift during fine-tuning or if includes search-generated data..
Learning to avoid and forget incorrect reasoning paths.
While NLL encourages the model to learn correct reasoning paths, it can only leverage the limited number of available successful trajectories and discards a large number of failed ones – which often vastly outnumber the correct ones (e.g., ranges from 5 to 400 in our experiments). To address the limitation of NLL, we propose leveraging failed trajectories by incorporating the unlikelihood (UL) loss (Welleck et al., 2019) as an additional objective. The UL loss enables the model to learn to avoid (i.e., forget) failed reasoning paths, effectively utilizing the large set of failed trajectories.
This objective explicitly reduces the probability of incorrect paths under the model. Unlike gradient ascent (GA) unlearning (Yao et al., 2023b)444The GA objective is unstable to optimize (Zhang et al., 2024b; d), as it has an unbounded optimum () and a divergent gradient near the optimum. , UL objective has a smooth optimization landscape, with gradient:
The gradient implies convergence to the stationary point where . From a REINFORCE view, Eq. 4 imposes an adaptive penalty of , penalizing higher-probability failures more strongly and thus efficiently suppressing wrong paths.
Combining learning and forgetting: unlikelihood fine-tuning (UFT).
Our final objective UFT combines NLL (Eq. 2) and UL losses (Eq. 3) with a coefficient :
Since correct paths provide explicit solutions while wrong paths only rule out alternatives, we treat UL as an auxiliary loss by setting close to 0. Moreover, as correct and failed paths may be drawn from the same search trees, they often share prefixes, introducing conflicts when is large, as also noted in preference-based methods (Zhang et al., 2024c). For example, suppose a prefix sequence () appears in both and with equal frequency. The gradient of Eq. 5 w.r.t. this prefix is , which has a stationary point at . Thus, sets the desired probability of the shared prefix, further motivating the need for a small to avoid conflicting objectives. Unlike prior work (Zhang et al., 2024c; Setlur et al., 2024) that identify high-credit steps with more computation, we sidestep this with an auxiliary loss without preference pairs.
4.3 Practical Algorithm
We summarize our method in Algo. 1. For a given reasoning task, we partition the initial state set into training, validation and test sets, namely, . Each subset has distinct input cases with follows the same instruction template. Our approach assumes access to a ground-truth verifier in (for generating training data) and (for checkpoint selection), but not in on which the final evaluation is conducted.
5 Experiments
We evaluate our approach on two mathematical reasoning tasks.
Game-of-24.
In Game-of-24555https://en.wikipedia.org/wiki/24_(puzzle), the goal is to use basic arithmetic operations (+-*/) and parentheses to combine four input numbers to obtain the number . Each input number can only be used once, which requires exactly three high-level reasoning steps. We follow the step-by-step response format from ToT (Yao et al., 2023a). For example, given numbers 2 9 10 12, a correct reasoning path is 12 * 2 = 24 (left: 9 10 24), 10 - 9 = 1 (left: 1 24), 24 * 1 = 24 (left: 24), Answer: (12 * 2) * (10 - 9) = 24. Each high-level step must include the remaining numbers after the operation. We implement a process-based verifier (Uesato et al., 2022) that checks both the intermediate steps and the final answer. For evaluation, we use the same test set from Yao et al. (2023a), consisting of cases. For training and validation, we use distinct cases, considered easier than the test set based on the human performance.
Countdown.
The Countdown game (Gandhi et al., 2024) extends Game-of-24 by requiring the use of four integers to obtain a specified target number beyond 24. All other rules remain the same including the step-by-step format. Following Gandhi et al. (2024), we randomly generate 500k training cases, excluding as a target number. For evaluation, we randomly generate 1,000 test cases with distinct target numbers from training set. This setup allows us to test generalization to Game-of-24 when trained on Countdown data.
| Training method | Inference method | Countdown test set | Game-of-24 test set | ||
| Succ rate | Time | Succ rate | Time | ||
| Qwen2.5-Math 7B (base model, no fine-tuning) | CoT (greedy) | 6.2% | 1.0 | 6% | 0.1 |
| CoT (pass@1) | 5.6% | 3.4 | 6% | 0.6 | |
| Search: ToT | 25% | 90 | 28% | 6.5 | |
| Search: RAP | 10.2% | 278 | 27% | 8.3 | |
| SFT baseline (, CoT data only) | CoT (greedy) | 32.5% | 0.5 | 16% | 0.1 |
| CoT (pass@1) | 32.2% | 3.1 | 12.3% | 0.6 | |
| Search: ToT | 22.2% | 96 | 30.7% | 8 | |
| Search: RAP | 17.5% | 141 | 27.7% | 16 | |
| Ours (, CoT or All data) | CoT (greedy) | 57.4% | 0.5 | 42% | 0.1 |
| CoT (pass@1) | 53.5% | 3.4 | 29.7% | 0.8 | |
| Search: ToT | 28.6% | 153 | 24% | 13 | |
| Search: RAP | 28.9% | 439 | 33.7% | 75 | |
5.1 Setup
Base LLMs.
We use Qwen2.5-Math 1.5B (Q1.5B) and Qwen2.5-Math 7B (Q7B) (Yang et al., 2024) as base LLMs. Qwen2.5-Math series were state-of-the-art open-weight mathematical LLMs as of December 2024.
Step 1: Data generation setup.
We mainly build upon the LLM-reasoners repository666https://github.com/maitrix-org/llm-reasoners (Hao et al., 2024), which provides a unified library for all the LLM reasoners studied here. We significantly accelerate inference by batching inputs using vLLM (Kwon et al., 2023). All reported inference time in our results is based on our optimized codebase. For each LLM reasoner algorithm, we vary its hyperparameters to gather data from . For CoT (Wei et al., 2022), we vary the temperature between and top-p between to have 9 variants. For each variant, we use a fixed 5-shot prompt template and collect paths for each case in Game-of-24 and paths for Countdown. For Game-of-24, we follow LLM-reasoners to implement ToT (Yao et al., 2023a) (varying the beam size from to ) and RAP (Hao et al., 2023) (varying the exploration parameter from to ). Since Countdown is implemented in Stream-of-Search (SoS) (Gandhi et al., 2024), we use SoS code to implement classic BFS and DFS algorithms, which do not use LLMs but rely on the oracle verifier. We merge the reasoning paths generated by all the variants of the same reasoner, and then remove duplicates by exact matching, to serve as our training dataset. See Appendix A for all details.
Step 2: Fine-tuning setup.
We follow the alignment handbook (Tunstall et al., 2024) to implement standard SFT and UFT on an instance with 8 A100 GPUs. We fix the batch size to and train epochs on each dataset for Game-of-24 and epochs for Countdown. We use a cosine schedule and sweep the peak learning rate over (1e-5, 5e-6, 2e-6) for Q1.5B and over (5e-6, 2e-6, 1e-6) for Q7B. We set to have standard SFT (NLL) and sweep over (1e-3, 1e-4, 1e-5, 1e-6) for UFT. See Appendix B for details.
Step 3: Evaluation on CoT and search-based reasoning.
Since our training dataset are all CoT-style reasoning paths, fine-tuning on these data should directly improve CoT reasoning capability. This is measured by performing zero-shot CoT (via greedy decoding or sampling) on with fine-tuned LLM. On the other hand, we also aim to retain search-based reasoning capability after fine-tuning. This is measured by performing ToT or RAP on .
Inference-time baselines on CoT and search-based reasoning.
Inference-time baselines refer to reasoners that use base LLMs without fine-tuning. For CoT reasoning, we use the few-shot CoT (via greedy-decoding or sampling) on the base LLMs. For search-based reasoning, we report the best variants of ToT and RAP with the base LLM. The results for these baselines are shown in Table 1 for Q7B and Table 5 for Q1.5B. We find that search-based reasoning (especially ToT) significantly improves success rates over CoT reasoning for Q7B (6.2% 25% in Countdown, 6% 28% in Game-of-24), but not for Q1.5B (2.2% 2.2% in Countdown, 2% to 8% in Game-of-24). However, search-based reasoning substantially increases inference time compared to CoT; for example, in Countdown, ToT and RAP are 90 and 278 slower than CoT (greedy decoding).
| Base | Countdown training data | Countdown test succ rate | |||||
| Sources | Quality | Base Inf | SFT (greedy) | UFT (greedy) | SFT (pass@1) | UFT (pass@1) | |
| Qwen2.5- Math 1.5B | CoT | 20.0% | 2.2% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.3% | ||||||
| Qwen2.5- Math 7B | CoT | 37.6% | 25% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.9% | ||||||
| Base | Countdown training data | Game-of-24 test succ rate | |||||
| Sources | Quality | Base Inf | SFT (greedy) | UFT (greedy) | SFT (pass@1) | UFT (pass@1) | |
| Qwen2.5- Math 1.5B | CoT | 20.0% | 8% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.3% | ||||||
| Qwen2.5- Math 7B | CoT | 37.6% | 28% | ||||
| BFS+DFS | 86.2% | ||||||
| CoT+BFS+DFS | 86.9% | ||||||
| Base | Game-of-24 training data | Game-of-24 test succ rate | |||||
| Sources | Quality | Base Inf | SFT (greedy) | UFT (greedy) | SFT (pass@1) | UFT (pass@1) | |
| Qwen2.5- Math 1.5B | CoT | 82.1% | 8% | ||||
| ToT+RAP | 44.7% | ||||||
| CoT+ToT+RAP | 87.6% | ||||||
| Qwen2.5- Math 7B | CoT | 92.9% | 28% | ||||
| ToT+RAP | 69.3% | ||||||
| CoT+ToT+RAP | 95.3% | ||||||
5.2 Boosting Chain-of-Thought Reasoning
We present CoT reasoning results at inference time for LLMs fine-tuned on the Countdown (2(b)) and the Game-of-24 datasets (Table 3). Similar to Ye et al. (2024), we also evaluate the models trained on Countdown on the Game-of-24 test set, because conceptually Game-of-24 is a subset of Countdown. For the Tables 2(b) and 3, we use the highest learning rate (1e-5 for Q1.5B and 5e-6 for Q7B) that yields best performance in CoT reasoning, leaving further discussion on the learning rate for Sec. 5.3. For the UFT objective, we fix the coefficient for the same model across different datasets, selected by the average success rate on validation set .
As noted in Sec. 4.2, prior work on SFT using synthetic data (Yuan et al., 2023; Singh et al., 2023; Yang et al., 2024) typically consider learning correct data generated by CoT only. Therefore, in both the Tables 2(b) and 3, only the gray cells (SFT on CoT data) represent the baselines, and all the other cells involve novelty of our approach (search data sources and/or UFT). We analyze these in detail below.
Training data quality is decisive to CoT reasoning.
First, we find that the test success rates are mainly determined by the training data quality, outweighing the effect of training objective (SFT or UFT). Data quality is measured by the best-of-n success rate: fraction of unique solved cases out of all the training cases.777Although ToT and RAP greatly outperform CoT as inference-time methods (Table 1, Table 5) in Game-of-24, they produce lower-quality training data than CoT (Table 3), because CoT’s best-of-n (n=100) success rate far exceeds its average (93% versus 6% in Q7B, 82% versus 5% in Q1.5B). The best performances in Table 3 and 2(b) correspond to the datasets with the highest quality (CoT data in Game-of-24, BFS+DFS data in Countdown). In addition, when the two datasets have similar quality (CoT vs CoT+ToT+RAP in Game-of-24, BFS+DFS vs CoT+BFS+DFS in Countdown), their performances are also alike. Remarkably, fine-tuning on the high-quality classic search (BFS+DFS) data from Countdown yields significantly better performance on Game-of-24 than models fine-tuned directly on Game-of-24 CoT data (37% vs 28.2% for Q7B and 29.7% vs 14.5% for Q1.5B). This highlights the importance of incorporating high-quality training data, rather than relying solely on CoT data, as is common practice.
Unlikelihood auxiliary loss helps CoT reasoning.
Next, we analyze the effect of incorporating UL loss (Eq. 5) to fine-tuning. By comparing SFT and UFT cells, we find UL helps improve performance in most scenarios: 14 out of 18 times for Q1.5B and 16 out of 18 for Q7B. While UFT provides a smaller boost than improving training data quality (typically under 4%), it still offers notable gains. For instance, UFT increases the success rate of Q7B fine-tuned on all data from 27.5% to 30.2%, surpassing its best inference-time baseline (28%). It also yields a substantial 7% improvement on the Game-of-24 test set for Q7B trained on all Countdown data.
5.3 Mitigating Forgetting in Search-based Reasoning
We investigate whether fine-tuned LLMs can retain search-based reasoning abilities (e.g., ToT, RAP), a question that remains underexplored in the absence of continued pretraining. Evaluating this retention is analogous to measuring backward transfer (Lopez-Paz & Ranzato, 2017) in continual learning, i.e., whether performance on a prior task (search-based reasoning) degrades after fine-tuning on a new one (CoT-style data). We focus on Q7B, which exhibits strong search-based reasoning performance and a clear gap over CoT baselines (Table 1). For completeness, the results of Q1.5B is shown in Appendix C.
Although learning rate decay is employed following the alignment handbook, we find that using a peak learning rate of 2e-5, as in the original Qwen2.5-Math setup (Yang et al., 2024), leads to catastrophic forgetting (McCloskey & Cohen, 1989), with near-zero performance in inference-time search. We hypothesize that this degradation stems from the lack of reward signals in CoT-style supervision, which weakens the model’s reward modeling ability essential for search-based reasoning (see Sec. A.4 for details).
A very low (peak) learning rate is crucial to retaining search capability.
Inspired by continual learning literature (Mirzadeh et al., 2020), we adopt a much smaller learning rate to prevent parameters from drifting too far from their initial state. Fig. 3 and Fig. 5 shows the effects of (peak) learning rate in standard SFT on Q7B. As we reduce the learning rate from 5e-6 to 1e-6 (an exceptionally low value for SFT), we observe a clear trade-off: CoT reasoning performance declines rapidly (though it remains above the baseline), while inference-time search improves significantly (though often remains below the baseline). To our knowledge, this is a novel finding in the LLM literature, where learning rate is rarely considered as a simple yet effective way to mitigate forgetting.
CoT training data is more effective for preserving search capability.
An interesting observation from Fig. 3 to Fig. 5 is that in both Countdown and Game-of-24 training settings, CoT data leads to better inference-time search than search-derived data, although CoT data has much lower quality compared to BFS+DFS data. Both CoT and search-derived data fall under CoT-style paths, but the latter is generated by a search algorithm rather than directly by the LLM. This suggests that CoT data, being more aligned with the LLM’s on-policy distribution, induces less distribution shift during SFT, thereby reducing forgetting.
Unlikelihood auxiliary loss can significantly improve search capability.
Building on the findings in Fig. 3, we observe that the effect of a low learning rate extends to UFT as well. Here, we examine the impact of the unlikelihood loss specifically at the lowest learning rate of 1e-6. Consistent with Sec. 5.2, UFT provides marginal improvements over SFT in CoT reasoning, regardless of the choice of . However, with =1e-4, inference-time search improves considerably compared to SFT (=0) in most cases. For example, with the same CoT data, ToT achieves 28.6% vs. 22.2% under SFT, and RAP reaches 28.9% vs. 17.5% – far exceeding the RAP baseline of 10.2%. This indicates that learning to forget incorrect CoT-style paths could generalize to forgetting incorrect intermediate search steps.
6 Conclusion
We propose a simple and effective method for mathematical reasoning, addressing three underexplored challenges: (1) unifying mixed-format data from diverse reasoners into a CoT-style format, (2) leveraging unpaired correct and failed paths via unlikelihood fine-tuning, and (3) mitigating loss of search capability using a small learning rate. Experiments on Countdown and Game-of-24 show that data quality, unlikelihood loss, and learning rate are key to balancing CoT efficiency and search capability.
Appendix A Reasoner Details
A.1 The Problems
This section provides details on the two math games not covered in the main paper.
Game-of-24.
We randomly split the first 900 cases (rank #1-900 by human average performance in the dataset888https://github.com/maitrix-org/llm-reasoners/blob/main/examples/ToT/game24/data/24.csv) into (720 cases) and (180 cases). The are the next 100 cases (rank #901-1000) following the setup in ToT (Yao et al., 2023a). As a result, the validation set evaluates in-distribution (ID) generalization, while the test set evaluates out-of-distribution (OOD) generalization based on human difficulty levels. Each case is guaranteed to be solvable using the four input numbers, which range from 1 to 13.
Countdown.
We follow SoS codebase999https://github.com/kanishkg/stream-of-search (Gandhi et al., 2024) to randomly generate 500k training cases, 1k validation cases, and 1k test cases. The training and validation sets share the same distribution of target numbers, while the test set includes distinct target numbers (including the number 24). Each problem consists of four input numbers sampled from 1 to 99, with a target number between 10 and 100. All generated problems are guaranteed to have at least one valid solution path using only integer intermediate results.
A.2 CoT Reasoner
Chain-of-Thought (Wei et al., 2022) uses the few-shot prompting template (see 1 for Countdown; Game-of-24 template follows the same format) for data generation and the zero-shot template (see 2) for fine-tuning and evaluation. The “greedy-decoding” evaluation uses zero temperature. The “pass@1” evaluation uses a temperature of 0.7 and top_p of 0.8 to sample 8 paths (n=8) for Countdown and 20 paths (n=20) for Game-of-24, then calculates the average success rate.
A.3 ToT and RAP Reasoners
The search-based reasoners (ToT (Yao et al., 2023a) and RAP (Hao et al., 2023)) are implemented using the LLM-reasoners codebase (Hao et al., 2024). Both use the same prompt templates but differ in their underlying search algorithm.
Tree-of-Thought.
ToT employs a beam search strategy. It begins with an empty set of nodes (the beam). At each search step, the LLM is prompted using the propose prompt (3) to generate plausible next steps for each node in the beam. Each proposed step is then evaluated with the intermediate reward prompt (4), which labels responses as “sure”, “likely”, or “impossible”. Following the original ToT setup, these labels are mapped to heuristic scores of 1, 0.1, and 0.0001, respectively. To reduce variance, we sample three responses for each reward step, and compute the average heuristic reward.
The beam is updated by selecting the top candidate nodes based on these heuristic scores, maintaining the predefined beam size. Since both Countdown and Game-of-24 tasks involve four steps in chain-of-thought reasoning, this process is repeated for four iterations. At the final search step, the terminal reward prompt (5) is used to assign a terminal reward to each node in the final beam. The path with the highest terminal reward is selected as the final reasoning path for evaluation, while all paths in the final beam are retained for training data.
To improve efficiency, we optimize the codebase using vLLM (Kwon et al., 2023), which enables batched inference across 1000 cases at the same time. In total, the ToT process requires passes (i.e., 8 calls to llm.generate() in vLLM) per batch. This batched approach significantly accelerates the search process, achieving over 100 speedup compared to the original for-loop execution in the LLM-Reasoners codebase.
In practice, we vary the beam size between 5 and 16 while keeping all other search parameters fixed during training data generation. Empirically, we find that a beam size of 5 for Countdown and 6 for Game-of-24 yields the best performance on Qwen2.5-Math 7B, while beam sizes of 8 and 10 perform best for the respective tasks on Qwen2.5-Math 1.5B. Accordingly, we adopt these optimal beam sizes for evaluating both the base LLMs and the improved models fine-tuned from them.
Reasoning-via-Planning.
RAP employs a Monte Carlo Tree Search (MCTS) strategy, maintaining a search tree initialized as empty. At each iteration, a node is selected for expansion using the Upper Confidence Bound applied to Trees (UCT) algorithm (Kocsis & Szepesvári, 2006), where the exploration parameter balances exploitation and exploration. The selected node is then expanded by simulating a full rollout: the propose prompt generates plausible next steps, and reward prompts are used to assign scores to each step along the way. Steps are sampled proportionally to their rewards throughout the rollout. After completing the rollout, the terminal reward is backpropagated through the tree to update the value estimates of the visited nodes.
We run this process for 100 iterations, resulting in up to passes. This is a loose upper bound, as many selected nodes are leaf nodes or near-terminal, leading to shorter rollouts in practice. While the original LLM-Reasoners implementation uses only 10 iterations, we increase it to 100 to improve performance, made feasible by our use of batched inference. At the end of the search, the path with the highest terminal reward is selected for evaluation, and all explored paths are extracted for training data.
In practice, we vary the exploration parameter over (1.0, 2.0, 4.0, 6.0, 8.0, 10.0) while keeping all other search parameters fixed during training data generation. Empirically, we find that an exploration parameter of 1.0 yields the best performance for both Countdown and Game-of-24 on Qwen2.5-Math 7B, whereas a value of 2.0 performs best for both tasks on Qwen2.5-Math 1.5B. Accordingly, we adopt these optimal exploration parameters for evaluating both the base LLMs and the improved models fine-tuned from them.
A.4 Loss of Search Capability after CoT Fine-Tuning: A Planning Perspective
An important insight emerges when comparing the prompting strategies used in CoT reasoning versus search-based methods like ToT and RAP: although they aim to solve the same task, they frame the role of the LLM in fundamentally different ways.
CoT treats the LLM as an open-loop controller, generating complete reasoning paths in a single pass (1) without any feedback or evaluation. In contrast, search-based methods integrates the LLM into a closed-loop system, where the LLM serves two interconnected roles: a policy that proposes next steps (3) and a reward model that evaluates them (4, 5). This interaction between proposal and intermediate feedback is central to structured planning and search.
This difference has meaningful consequences for fine-tuning. When an LLM is fine-tuned on CoT-style paths that strip away reward signals, it may weaken its ability to perform reward modeling, which is critical in search-based reasoning. This observation offers a planning-theoretic explanation for why models fine-tuned on CoT-style paths may struggle with inference-time search.
A.5 Classic BFS and DFS Reasoners
We also employ classic search algorithms, Breadth-First Search (BFS) and Depth-First Search (DFS), to generate training data for the Countdown task. We adopt the implementation from SoS (Gandhi et al., 2024) without modification. These symbolic reasoners are significantly faster than LLM reasoners and achieve high success rates of 65.4% (BFS) and 81.5% (DFS) on the training set (due to the use of pruning heuristics, they do not guarantee a solution in all cases). However, unlike LLM reasoners, these methods rely on external tools such as calculators and goal checkers to guide the search process.
Appendix B Fine-tuning Details
We follow the alignment handbook101010https://github.com/huggingface/alignment-handbook (Tunstall et al., 2024) to fine-tune two base LLMs, Qwen2.5-Math 1.5B (Q1.5B)111111https://huggingface.co/Qwen/Qwen2.5-Math-1.5B and Qwen2.5-Math 7B (Q7B)121212https://huggingface.co/Qwen/Qwen2.5-Math-7B (Yang et al., 2024).
We extend the training codebase to support Unlikelihood Fine-Tuning (UFT), where UFT and Supervised Fine-Tuning (SFT) differ only by the unlikelihood loss coefficient . A batch size is used for both successful and failed data batches (after gradient accumulation). We disable sequence packing to prevent cross-contamination131313https://github.com/huggingface/transformers/issues/25452 and set a context length of 256 tokens. which is sufficient for our tasks. Training is performed in bfloat16 precision, and model checkpoints are saved every 5% epochs for later evaluation.
The learning rate follows cosine_with_min_lr schedule141414https://github.com/huggingface/transformers/blob/a22a4378d97d06b7a1d9abad6e0086d30fdea199/src/transformers/optimization.py#L338C5-L338C48: it linearly warms up from 0 to over the first 10% epochs, then decays to a minimum of 7e-8 via cosine function for the rest epochs. In Sec. 5.2, we use a peak learning rate of 1e-5 for Q1.5B and 5e-6 for Q7B, while in Sec. 5.3, we decrease to much lower values, as discussed in the main paper.
The training time is measured in number of epochs based on the correct dataset size . Although training runs for multiple epochs, each failed path is shown to the model only once, as the pool of failed examples is significantly larger than the correct dataset.
For Countdown, BFS+DFS generate 614k correct paths. The full CoT+BFS+DFS dataset contains 734k correct paths for Q1.5B and 985k for Q7B. Due to the large dataset size, we train for 2 epochs. The training time on CoT+BFS+DFS data is less than 5 hrs for Q1.5B and 30 hrs for Q7B on an instance with 8 A100s (40 GB). For Game-of-24, CoT+ToT+RAP contains 6k correct paths for Q1.5B and 14k for Q7B. We train for 10 epochs, with total training time under 0.5 hrs for Q1.5B and 2.5 hrs for Q7B.
We emphasize that our goal is to reduce inference time in reasoning. While training may be relatively expensive, it is performed offline and does not impact deployment efficiency or user-facing latency.
Appendix C Additional Results
Limited effects on 1.5B due to weak base performance.
Although not the primary focus of this work, we provide results for Q1.5B in Appendix (Fig. 6 to Fig. 8) for completeness. Given its weak search-based reasoning performance as a base model (Table 5; success rate 2% in Countdown), the effects of learning rate, CoT data, and unlikelihood loss are minimal, likely due to its poor initial capability. This suggests that the observed benefits may be more pronounced at larger model scales.
Appendix D Limitation
This paper focuses on improving mathematical-specific LLMs. Preliminary experiments on generic models like Mistral 7B (Jiang et al., 2023) show that our pipeline remains effective, providing similar benefits as on Qwen2.5-Math 1.5B, except that unlikelihood loss offers little improvement over NLL on the same training data.
| Training method | Inference method | Countdown test set | Game-of-24 test set | ||
| Succ rate | Time | Succ rate | Time | ||
| Qwen2.5-Math 7B (base model, no fine-tuning) | CoT (greedy) | 6.2% | 1.0 | 6% | 0.1 |
| CoT (pass@1) | 5.6% | 3.4 | 6% | 0.6 | |
| search: ToT | 25% | 90 | 28% | 6.5 | |
| search: RAP | 10.2% | 278 | 27% | 8.3 | |
| SFT baseline (lr=5e-6, =0, CoT data) | CoT (greedy) | 32.5% | 0.5 | 16% | 0.1 |
| CoT (pass@1) | 32.2% | 3.1 | 12.3% | 0.6 | |
| search: ToT | 4% | 109 | 3% | 7.9 | |
| search: RAP | 5.7% | 146 | 1% | 14 | |
| SFT baseline (lr=1e-6, =0, CoT data) | CoT (greedy) | 24.7% | 0.5 | 8% | 0.1 |
| CoT (pass@1) | 24.1% | 2.9 | 7.3% | 0.6 | |
| search: ToT | 22.2% | 96 | 30.7% | 8 | |
| search: RAP | 17.5% | 141 | 27.7% | 16 | |
| Ours (lr=5e-6, =1e-3, All data) | CoT (greedy) | 57.4% | 0.5 | 42% | 0.1 |
| CoT (pass@1) | 53.5% | 3.4 | 29.7% | 0.8 | |
| search: ToT | 0% | 1.9 | 0.3% | 0.4 | |
| search: RAP | 0% | 341 | 3.7% | 29 | |
| Ours (lr=1e-6, =1e-4, CoT data) | CoT (greedy) | 24.7% | 0.5 | 8.3% | 0.1 |
| CoT (pass@1) | 24.7% | 3.3 | 6.7% | 0.8 | |
| search: ToT | 28.6% | 153 | 24% | 13 | |
| search: RAP | 28.9% | 439 | 33.7% | 75 | |
| Training method | Inference method | Countdown test set | Game-of-24 test set | ||
| Succ rate | Time | Succ rate | Time | ||
| Qwen2.5-Math 1.5B | CoT (greedy) | 2.2% | 0.6 | 2% | 0.1 |
| CoT (pass@1) | 2.1% | 3.2 | 3% | 0.5 | |
| search: ToT | 2.2% | 68 | 8% | 5.4 | |
| search: RAP | 1.4% | 143 | 4% | 11.3 | |
| SFT baseline (lr=1e-5, =0, CoT data) | CoT (greedy) | 24.2% | 0.5 | 13.3% | 0.1 |
| CoT (pass@1) | 24.4% | 2.8 | 9.3% | 0.5 | |
| search: ToT | 1.1% | 85 | 1.7% | 3.1 | |
| search: RAP | 3.3% | 122 | 3.7% | 7.3 | |
| Ours (lr=1e-5, =1e-3, All data) | CoT (greedy) | 51.5% | 0.3 | 26% | 0.1 |
| CoT (pass@1) | 46.5% | 2.3 | 25% | 0.5 | |
| search: ToT | 0% | 1.0 | 0% | 0.1 | |
| search: RAP | 0% | 88 | 0% | 9.5 | |